there are 12 more history books than math books in the library, if there are 122 sience and math book how many math books are there in the library.
Answer by checkley77(12569)

(12+M)+M=122
12+2M=122
2M=122-12
M=110/2
M=55 NUMBER OF MATH BOOKS.
PROOF:
12+55+55=122
122=122